On March 31, 2021, Top Notch Goods committed to a plan to sell machinery equipment. The machinery equipment was available for immediate sale, and an active plan to locate a buyer was initiated. The machinery equipment had been purchased on January 1, 2019, for $260,000. The machinery equipment had an estimated six-year service life and residual value of $20,000. The machinery equipment was being depreciated using the straight-line method. Top Notch Good's fiscal year ends on December 31.

  1. Calculate the equipment’s book value as of March 31, 2021

  2. By December 31, 2021, the machinery equipment has not been sold, but management expects that it will be sold in 2022 for $150,000. For what amount is the machinery equipment reported in the December 31, 2021, balance sheet?
Answer #1

a) Equipment Book value = Cost-Accumulated depreciation = 260000-(240000/72*27) = 170000

b) Equipment book value = Cost -accumulated depreciation = 260000-(240000/6*3) = 140000